# A splitting method for numerical relativistic magnetohydrodynamics

### v1

@@ 1 Introduction
The strong gravity of astrophysical black holes and neutron stars creates some of the most extreme physical conditions in the Universe which cannot be achieved in research laboratories. In particular, they naturally develop magnetospheres with extremely high plasma magnetisation. Highly-relativistic winds and jets emerging from these magnetospheres create spectacular structures on enormous scales, from parsecs (pulsar wind nebulae) to hundreds of kiloparsecs (jets of active galactic nuclei). These flows transport huge amounts of energy in the form of Poynting flux and the kinetic energy of the bulk motion, and drive the observed phenomena via magnetic reconnection and shock interaction with the external plasma. For plasma flows on such huge scales, relativistic magnetohydrodynamics (RMHD) is the most suitable framework. In contrast to the non-relativistic problems, where the magnetisation is well described by the ratio of thermal and magnetic pressures ($\beta=p/p_{m}$), the magnetisation of relativistic plasma is best described by the parameter $\sigma=\hat{B}^{2}/4\pi w$, where $\hat{B}$ is the magnetic field strength measured in the rest frame of plasma and $w$ is the relativistic enthalpy of plasma which includes the contribution due to the rest mass energy of plasma particles.
The most popular numerical schemes for compressible astrophysical fluid dynamics and MHD at present time are fully conservative. One of the key advantages of such schemes is their ability to capture shock waves very accurately. Following the trend, a number of computer codes for RMHD based on this approach appeared over the last two decades, starting from Komissarov (1999), and they have been very useful in many applications. However, almost from the start it also became apparent that this approach has a severe limitation - the conservative codes have the tendency to crash in problems involving highly-magnetised regions as the updated set of conserved variables (like the total energy and momentum) could not be converted into physically meaningful set of primitive variable (like the thermal pressure and velocity of plasma). In multi-dimensional simulations, these schemes begin to fail when $\sigma\sim 1$. On the one hand, this is a very high magnetisation, never achieved in laboratory plasmas. On the other hand, it can be much higher in many problems of relativistic astrophysics. For example, in the magnetospheres of neutron stars, $\sigma$ can be as high as $10^{3}-10^{6}$.
For adiabatic flows, one can eliminate the energy equation from the set of numerically integrated equations of RMHD (Komissarov et al., 2007b; Noble et al., 2009), which helps to extend the range of manageable magnetisation up to $\sigma\approx 100$. The conversion of remaining conserved variables to the primitive variables may still fail from time to time, becoming increasingly more severe for higher $\sigma$ and requiring emergency fixes just to keep the simulations going. Most importantly, however, this approach severely limits the range of applications, excluding the problems involving shocks and current sheets.
Excessive artificial plasma heating due to numerical resistivity is another issue in simulations of highly magnetised flows with standard conservative schemes for RMHD. For this reason, the high-$\sigma$ region is normally excluded when modelling black hole emission in simulations (e.g., Event Horizon Telescope Collaboration et al., 2021).
@@
It has been suggested that the origin of this issue is the stiffness of the conservation laws of RMHD in the high-$\sigma$ regime (Komissarov, 2006a). Basically, when $\sigma$ is high, the electromagnetic field dominates in the total energy and momentum. In this case, it is reasonable to expect that even small errors in the magnetic field, associated with the numerical integration of the Faraday equation, may result in large errors in the plasma parameters, when they are computed from the conserved quantities. The quantitative analysis of the errors is rather complicated, however. To strengthen the argument, one may approach this problem from a different angle. The dynamics of electromagnetic field in highly rarefied plasma can be described using the approximation of the Force-Free Degenerate Electrodynamics (FFDE, e.g. MacDonald & Thorne, 1982; Uchida, 1997; Gruzinov, 1999; Komissarov, 2002)). Normally, it is formulated as the Maxwell equations complimented with a constraint on the electric 4-current, which ensures vanishing Lorentz-force. The density of electric charges required to satisfy this constraint is quite small and the corresponding energy-momentum density of plasma can be negligibly small compared to that of the electromagnetic field. Alternatively, one may consider FFDE as RMHD in the limit $\sigma\to\infty$ (Komissarov, 2002). In this limit, the set of the differential equations of RMHD reduces to the Faraday equation and the energy-momentum conservation laws for the electromagnetic field, complemented with the two perfect conductivity conditions. However, this system is overdetermined, with only two out of the four components of the energy-momentum equation being independent. For the numerical integration, this implies that any error in the computed magnetic field makes it inconsistent with the computed energy-momentum density. The omission of the energy equation in the simulations of adiabatic flows reduces the stiffness, which explains why this allows us to deal with higher plasma magnetisation.
@@
A similar difficulty was identified some time ago in the MHD simulations of the collision between the Earth’s magnetosphere and the solar wind. In this problem, the Earth’s magnetic field increases by many orders of magnitude from the collision site to the troposphere, where it is largely stationary and dipolar, whereas the perturbation of this field remains of about the same magnitude and hence increasingly small relative to the undisturbed Earth’s field on approach to the troposphere. If the total magnetic field is evolved numerically, the truncation errors become large in comparison to the perturbation amplitude and hence the numerical solution for the perturbation becomes corrupted. To overcome this problem, Tanaka (1994) proposed to separate the strong stationary dipolar field from its perturbation and hence to integrate only the nonlinear equations governing the perturbation. This approach has been very effective and it is now standard in numerical modelling of planetary magnetospheres (e.g. Eggington et al., 2020).
@@
Our problem is more complicated, as in the most interesting astrophysical applications, the strong background magnetic field is highly dynamic and cannot be considered as a known stationary component. At the first sight, this could be handled by allowing it to evolve according to the evolution equations of FFDE. However, over time, the RMHD solution for the electromagnetic field could significantly deviate from the FFDE solution, with the force-free component and its perturbation having similar amplitudes. To keep the electromagnetic perturbation small, one could reset the division of the electromagnetic field into the strong force-free and perturbation components. The simplest way of doing this is to recombine the force-free field and its perturbation into the ’refreshed’ force-free field, and to nullify the perturbation at the same time. In sufficiently simple problems, this could be done only so often. However, in some other problems, the perturbation may grow very rapidly. For example, consider a stationary fast magnetosonic shock. Since the fast modes of FFDE propagate with the speed of light, the FFDE solution will strongly deviate from the RMHD solution already after one time step of numerical integration. This shows that to make the scheme robust, one has to invoke the resetting every time step.
@@
Numerical integration of FFDE equations, either in the form of the Maxwell equations with force-free current (Gruzinov, 1999) or in the form of reduced RMHD equations (Komissarov, 2002), cannot preserved full conservation of electromagnetic energy-momentum and hence the total energy-momentum. Thus, the splitting approach constitutes a major departure from modern conservative schemes for RMHD. The departure seems inevitable for high-$\sigma$ RMHD as it is the attempt to ensure the full conservation that leads to the crashes.
@@
In this paper, we describe a successful attempt to develop the splitting method based on these ideas. In section 2, we detail the key principles of the splitting method. Section 3 describes the specifics of its numerical implementation. The 1D test simulations are presented in section 4. In addition to the standard tests involving hyperbolic waves of RMHD, this section also describes the investigation of the scheme’s numerical resistivity and the possibility to control the plasma heating via numerical magnetic dissipation. Section 5 describes the test simulations for inherently 2D problems. These include the investigation of the anisotropy of numerical resistivity, and a number of problems involving current sheets. The latter constitute a study focusing on the ability of ideal MHD codes to capture fast magnetic reconnection, apparently the first study of this kind. The whole study is summarised in section 6 and the key conclusions are stated in section 7. Appendix A provides with the derivation of the key equations involved in the variable conversion algorithm, and appendix B gives the exact solutions of shock equations, used in the test simulations of magnetohydrodynamic shocks.
@@ 2.1 Ideal Relativistic Magnetohydrodynamics
The 3+1 system of ideal RMHD in Minkowski spacetime consists of the Faraday equation
@@
$\partial_{t}\left(\frac{E^{2}+B^{2}}{2}+w\gamma^{2}-p\right)+\nabla\!\cdot\!(\mn@boldsymbol{E}\!\times\!\mn@boldsymbol{B}+w\gamma^{2}\mn@boldsymbol{v})=0\,,$ (2)
@@
$\partial_{t}\left(\mn@boldsymbol{E}\!\times\!\mn@boldsymbol{B}+w\gamma^{2}\mn@boldsymbol{v}\right)+\nabla\!\cdot\!\left[-\mn@boldsymbol{E}\otimes\mn@boldsymbol{E}-\mn@boldsymbol{B}\otimes\mn@boldsymbol{B}+w\gamma^{2}\mn@boldsymbol{v}\otimes\mn@boldsymbol{v}+\mbox{\bf{g}}\left(\frac{E^{2}+B^{2}}{2}+p\right)\right]=0\,,$ (3)
@@
$\partial_{t}(\rho\gamma)+\nabla\!\cdot\!(\rho\gamma\mn@boldsymbol{v})=0\,,$ (4)
@@
Here $p$ is the thermodynamic pressure, $\rho$ is the density of plasma particles rest mass, ${\mathbf{g}}$ is the metric tensor of space, ${v}$ is the fluid velocity, $\gamma$ is the corresponding Lorentz factor, ${B}$ and ${E}$ are the vectors of electric and magnetic field respectively. $w(p,\rho)$ is the relativistic enthalpy per unit volume. In what follows, we use the equation of state
@@
where $\Gamma$ is the ratio of specific heats. Here we utilise the relativistic units where neither the speed of light no the geometric factor $1/4\pi$ appears in the equations. We also agree that for any 3-vector of the space ${a}$, $a^{2}=a_{i}a^{i}$ and $a=\sqrt{a^{2}}$, and for any 4-vector $a_{\nu}$ of the spacetime, $a^{2}=a_{\nu}a^{\nu}$.
Let us now analyse the energy-momentum conservation to see how its numerical integration can result in an unphysical state. Consider the 4-vector of energy-momentum density, $\Pi^{\mu}=T^{\mu\nu}n_{\nu}$, where $T^{\mu\nu}$ is stress-energy-momentum tensor, and $n_{\nu}$ is the 4-velocity of the fiducial observer who measures the energy and momentum. When the observer is at rest in the space, $n_{\nu}=\delta^{t}_{\nu}$ and $\Pi^{\nu}=({\cal E},\mn@boldsymbol{S})$, where ${\cal E}$ and ${S}$ are the energy and momentum densities respectively. For the electromagnetic field, this is
$\Pi_{\mbox{\tiny em}}^{\nu}=\left(\frac{E^{2}+B^{2}}{2},\mn@boldsymbol{E}\!\times\!\mn@boldsymbol{B}\right)\,,\quad\mbox{where}\quad\mn@boldsymbol{E}=-\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}\,,$
@@
$\Pi_{\mbox{\tiny em}}^{2}=-\frac{1}{4}\left(B_{\parallel}^{2}+\frac{B_{\perp}^{2}}{\gamma^{2}}\right)^{2}<0\,,$
@@
$\Pi^{\nu}_{\mbox{\tiny pl}}=(w\gamma^{2}-p,w\gamma^{2}\mn@boldsymbol{v})\,,$
@@
$\Pi_{\mbox{\tiny pl}}^{2}=\gamma^{2}w(2p-w)-p^{2}\,.$
@@
$\Pi_{\mbox{\tiny t}}^{2}=\Pi_{\mbox{\tiny em}}^{2}+\Pi_{\mbox{\tiny pl}}^{2}+2(-{\cal E}_{\mbox{\tiny em}}{\cal E}_{\mbox{\tiny pl}}+(\mn@boldsymbol{S}_{\mbox{\tiny em}}\!\cdot\!\mn@boldsymbol{S}_{\mbox{\tiny pl}}))<0$

### v3

@@ 1 Introduction
The strong gravity of astrophysical black holes and neutron stars creates some of the most extreme physical conditions in the Universe which cannot be achieved in research laboratories. In particular, they naturally develop magnetospheres with extremely high plasma magnetisation. Highly-relativistic winds and jets emerging from these magnetospheres create spectacular structures on enormous scales, from parsecs (pulsar wind nebulae) to hundreds of kiloparsecs (jets of active galactic nuclei). These flows transport huge amounts of energy in the form of Poynting flux and the kinetic energy of the bulk motion, and drive the observed phenomena via magnetic reconnection and shock interaction with the external plasma. For plasma flows on such huge scales, relativistic magnetohydrodynamics (RMHD) is the most suitable framework.
The starting point of modern numerical schemes for compressible hydrodynamics (HD) and magnetohydrodynamics (MHD) is their differential equations written in the form of conservation laws, with the aim of developing a numerical analog of these laws which provides conservation of the scheme-specific numerical approximation for the conserved quantities integrated over the volume of individual computational cells and the whole computational domain down to the processor rounding (machine) error. This is mainly because of the superior ability of such schemes to accurately capture shock waves. Both finite-volume and finite-difference schemes can be developed along these lines. The schemes may differ in many other aspects, like the type of Riemann solvers, the order of accuracy, etc. In terms of the integration of the Faraday equation, the modern numerical schemes for MHD and RMHD split into two major groups, those which use the generalised method of Lagrange multiplier (GLM, Munz et al., 2000; Dedner et al., 2002) and those that use the constrained transport method (CT, Evans & Hawley, 1988).
In form, the Faraday equation of the MHD (and RMHD) system can be written as a law describing the conservation of the magnetic field integrated over the volume. This approach allows to integrate the Faraday equation in the same fashion as the other conservation laws, which is very convenient. However, it makes impossible to ensure that the magnetic field remains divergence-free, whatever numerical approximation of the divergence is chosen, because this condition involves not the volume integral of the magnetic field but its integral (magnetic flux) over a close surface. The uncontrolled deviation from the divergence-free state is not just an error for the magnetic field alone but may have severe implication for other flow parameters (Brackbill & Barnes, 1980). For this reason, additional algorithms were developed to keep the magnetic field close to a divergence-free state.
In the CT method, the Faraday equation is treated as the law describing the conservation of magnetic flux over a surface. As the result, the normal components of the magnetic field have to be defined at the faces of computational cells and the electric field over their edges. This leads to the so-called staggered grid, where different physical parameters are defined at different points of the grid making this approach rather involved (for the recent analysis and comparison of various CT schemes see Mignone & Del Zanna, 2021). Moreover, since the energy (in Newtonian MHD) and the whole energy-momentum vector (in RMHD) include the electromagnetic contribution, integration of the corresponding conservation laws requires to define the magnetic field inside the computational cells as well. As a result, there are two numerical solutions for the magnetic fields, one defined at the cell faces and advanced using the surface form of the integral Faraday equation, and one defined inside the cells and advanced using the volume form of this equation. They may diverge over time, resulting in a rather uncomfortable state where the volume-based conserved variables like energy and momentum are inconsistent with the face-based magnetic field. In highly magnetised regions, this may even lead to an unphysical state with negative gas pressure. To prevent this, one may readjust the volume-based magnetic field and hence the conserved variables, thus making the numerical scheme not fully conservative. In Newtonian MHD, where out of all other conserved quantities the magnetic field is present only in the total energy, this can be done by simply removing the energy contribution due to the volume-based magnetic field and replacing it with the contribution due to the surface-base magnetic field (Balsara & Spicer, 1999). This algorithm keeps the plasma contribution to the energy and hence the corresponding primitive variables (velocity, density and thermal pressure of plasma) unchanged, thus making the deviation from the energy conservation superficial. In RMHD, both the magnetic and the electric field contribute to all the components of the total energy-momentum vector. Even if the volume-based magnetic field is a conserved variable, the volume-based electric field is not and can be found only via conversion of the conserved variables into the primitive ones (velocity, density and thermal pressure of plasma) first, and then applying the perfect conductivity condition ${E}=-\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}$. As a result, the direct replacement of the electromagnetic contribution to the energy-momentum like in Balsara & Spicer (1999) is impossible. Instead, one is forced to carry out the “harmonisation” only after the variable conversion (Komissarov, 1999). The same algorithm can be applied in Newtonian MHD, where it is equivalent to that of Balsara & Spicer (1999).
In the GLM method, the system of differential equations is extended by 1) introducing new scalar dependent variable and hence one more evolution equation, and 2) modifying the Faraday equation to couple the new variable and the magnetic field. As a result, ${\nabla}\!\cdot\!{B}$ can be both transported away from the regions where it is generated, normally regions with large computational errors, and diffused over the computational domain effectively dumping its spacial oscillation. However, some residual divergence of the magnetic field remains. This method is very easy to implement, as it does not require any modification of the computational grid, and all the evolution equations are treated in the same way. At least in ideal Newtonian MHD, the GLM and CT methods provide similar accuracy (Mignone et al., 2010).
In contrast to Newtonian MHD, where it is sufficient to describe the plasma magnetisation by the ratio of thermal and magnetic pressures ($\beta=p/p_{\mbox{\tiny m}}$), relativistic plasma requires a different parameter $\sigma=b^{2}/4\pi w$, where $w=\epsilon+p$ is the relativistic enthalpy of plasma (where the internal energy density $\epsilon$ includes the contribution due to the rest mass energy of plasma particles), and $b^{2}=B^{2}-E^{2}$ is the Lorentz invariant equal to the squared strength of magnetic field as measured in the rest frame of plasma. Direct extension of the numerical methods developed for Newtonian MHD to RMHD has been quite successful in the low-$\sigma$ regime. However, the high-$\sigma$ regime has turned out to be very problematic, as these codes tend to crash, resulting in conserved variables which cannot be converted into physically meaningful set of primitive variables. In multi-dimensional simulations, these schemes begin to fail when $\sigma\sim 1$. On the one hand, this is a very high magnetisation, never achieved in laboratory plasmas. On the other hand, it can be much higher in many problems of relativistic astrophysics. For example, in the pulsar magnetospheres, $\sigma$ can be as high as $10^{3}-10^{6}$.
@@
It has been suggested that the origin of this issue is the stiffness of the conservation laws of RMHD in the high-$\sigma$ regime (Komissarov, 2006a). Basically, when $\sigma$ is high, the electromagnetic field dominates in the total energy and momentum. In this case, it is reasonable to expect that even small errors in the magnetic field, associated with the numerical integration of the Faraday equation, may lead to large errors in the plasma parameters, when they are computed from the conserved quantities. The quantitative analysis of the errors is rather complicated, however. To strengthen the argument, one may approach this problem from a different angle. The dynamics of electromagnetic field in highly rarefied plasma can be described using the approximation of the Force-Free Degenerate Electrodynamics (FFDE, e.g. MacDonald & Thorne, 1982; Uchida, 1997; Gruzinov, 1999; Komissarov, 2002)). Normally, it is formulated as the Maxwell equations complimented with a constraint on the electric 4-current, which ensures vanishing Lorentz-force. The density of electric charges required to satisfy this constraint is quite small and the corresponding energy-momentum density of plasma can be negligibly small compared to that of the electromagnetic field. Alternatively, one may consider FFDE as RMHD in the limit $\sigma\to\infty$ (Komissarov, 2002). In this limit, the set of the differential equations of RMHD reduces to the Faraday equation and the energy-momentum conservation laws for the electromagnetic field, complemented with the two perfect conductivity conditions. However, this system is overdetermined, with only two out of the four components of the energy-momentum equation being independent. For the numerical integration, this implies that any error in the computed magnetic field makes it inconsistent with the computed energy-momentum density.
For adiabatic flows, one can eliminate the energy equation from the set of numerically integrated equations of RMHD and this helps to extend the range of manageable magnetisation up to $\sigma\approx 100$ (Komissarov et al., 2007b; Noble et al., 2009). The conversion of remaining conserved variables to the primitive variables may still fail from time to time, becoming increasingly more severe for higher $\sigma$ and requiring emergency fixes just to keep the simulations going. Nonetheless, these results support the stiffness of RMHD equations as the reason for the high-$\sigma$ failures, as the omission of the energy equation reduces this stiffness.
In their CT scheme for ideal RMHD, Mignone & Bodo (2006) applied exactly the same energy correction as Balsara & Spicer (1999), before converting the conserved variables into the primitive ones and replacing the cell-based magnetic field with the face-based one. This allowed them to succeed with the cylindrical explosion test for $B_{0}=1$ (see section 5.3), which would failed otherwise. Martí (2015) went further and proposed an iterative algorithm for correcting the total energy-momentum vector, using the correction of Mignone & Bodo (2006) as a first step. This allowed them to avoid the conversion failure in the explosion test with $B_{0}=100$. The corresponding plasma magnetisation in this problem is extremely high, with $\sigma>2.5\times 10^{3}$! This is a remarkable achievement, but some caution is in order. First, these corrections of conserved variables are different to that of Balsara & Spicer (1999) as they change not only the energy-momentum of the electromagnetic field but the energy-momentum of plasma as well. So, this is more than just resetting the cell-based magnetic field using the face-centred magnetic field. Second, there is no analysis to justify these corrections. The mathematical problem supposed to be solved by the iterative algorithm of Martí (2015) is actually not posed. Even a qualitative understanding of why these corrections help to keep the conserved variables in the physical part of their domain is missing. As the result, it is not clear how robust these fixes are. A proper analytical and numerical investigation of this approach is required.
In addition to the conversion failures, RMHD simulations of highly magnetised flows suffer from excessive artificial plasma heating due to numerical resistivity. For this reason, the high-$\sigma$ region is normally excluded when modelling black hole emission in simulations (e.g., Event Horizon Telescope Collaboration et al., 2021).
@@
In this paper, we propose an alternative to emergency fixes in the form of a radically new approach to computational RMHD which allows to mitigate the stiffness of its equations in the high-$\sigma$ limit. In part, it was motivated by the method used by Tanaka (1994) for MHD simulations of the collision between the Earth’s magnetosphere and the solar wind. In this problem, the Earth’s magnetic field increases by many orders of magnitude from the collision site to the troposphere, where it is largely stationary and dipolar, whereas the perturbation of this field remains of about the same magnitude and hence increasingly small relative to the undisturbed Earth’s field on approach to the troposphere. If the total magnetic field is evolved numerically, the truncation errors become large in comparison to the perturbation amplitude and hence the numerical solution for the perturbation becomes corrupted. To overcome this problem, Tanaka (1994) proposed to separate the strong stationary dipolar field from its perturbation and hence to integrate only the nonlinear equations governing the perturbation. This approach has been proved to be very effective, and nowadays it is widely used in numerical modelling of planetary magnetospheres (e.g. Eggington et al., 2020).
@@
Our problem is more complicated, however, as in the most interesting astrophysical applications, the strong background magnetic field is highly dynamic and cannot be considered as a known stationary component. At the first sight, this could be handled by allowing it to evolve according to the evolution equations of FFDE (Komissarov, 2001, 2006b; McKinney, 2006; Mahlmann et al., 2021). However, over time, the RMHD solution for the electromagnetic field could significantly deviate from the FFDE solution, with the force-free component and its perturbation having similar amplitudes. To keep the electromagnetic perturbation small, one could reset the division of the electromagnetic field into the strong force-free and perturbation components. The simplest way of doing this is to recombine the force-free field and its perturbation into the ’refreshed’ force-free field, and to nullify the perturbation at the same time. In sufficiently simple problems, this could be done only so often. However, in some other problems, the perturbation may grow very rapidly. For example, consider a stationary fast magnetosonic shock. Since the fast modes of FFDE propagate with the speed of light, the FFDE solution will strongly deviate from the RMHD solution already after one time step of numerical integration. This shows that to make the scheme robust, one has to invoke the resetting every time step.
@@
Numerical integration of FFDE equations, either in the form of the Maxwell equations with force-free current (Gruzinov, 1999) or in the form of reduced RMHD equations (Komissarov, 2002), does not conserve the electromagnetic energy-momentum and hence the splitting approach cannot ensure the conservation of the total energy-momentum down to the processor rounding error. However, a departure from this conservation seems inevitable for high-$\sigma$ RMHD anyway, because it is the attempt to ensure the full conservation that leads to crashes.
@@
In this paper, we describe a successful attempt to develop the splitting method based on these ideas. In section 2, we detail the key principles of the splitting method. Section 3 describes the specifics of its numerical implementation. The 1D test simulations are presented in section 4. In addition to the standard tests involving hyperbolic waves of RMHD, this section also describes the investigation of the scheme’s numerical resistivity and the possibility to control the plasma heating via numerical magnetic dissipation. Section 5 describes the test simulations for inherently 2D problems. These include the investigation of the anisotropy of numerical resistivity, and a number of problems involving current sheets. The latter constitute a study focusing on the ability of ideal MHD codes to capture fast magnetic reconnection, apparently the first study of this kind. The whole study is summarised in section 6 and the key conclusions are stated in section 7. Appendix A describes the novel 3rd-order WENO interpolation employed by our code, and Appendix B gives the derivation of the key equations involved in the variable conversion algorithm.
@@ 2.1 Ideal Relativistic Magnetohydrodynamics
For an inertial frame of Minkowski spacetime, the system of ideal RMHD in consists of the Faraday equation
@@
$\partial_{t}\left(\frac{E^{2}+B^{2}}{2}+w\gamma^{2}-p\right)+\mn@boldsymbol{\nabla}\!\cdot\!(\mn@boldsymbol{E}\!\times\!\mn@boldsymbol{B}+w\gamma^{2}\mn@boldsymbol{v})=0\,,$ (2)
@@
$\partial_{t}\left(\mn@boldsymbol{E}\!\times\!\mn@boldsymbol{B}+w\gamma^{2}\mn@boldsymbol{v}\right)+$
[${E}\otimes\mn@boldsymbol{E}-\mn@boldsymbol{B}\otimes\mn@boldsymbol{B}+w\gamma^{2}\mn@boldsymbol{v}\otimes\mn@boldsymbol{v}+\mbox{\bf{g}}\left(\frac{E^{2}+B^{2}}{2}+p\right)\right]=0\,,$ (3)
@@
$\partial_{t}(\rho\gamma)+\mn@boldsymbol{\nabla}\!\cdot\!(\rho\gamma\mn@boldsymbol{v})=0\,,$ (4)
@@
Here $p$ is the thermodynamic pressure, $\rho$ is the density of plasma particles rest mass, ${\mathbf{g}}$ is the metric tensor of space, ${v}$ is the fluid velocity, $\gamma$ is the corresponding Lorentz factor, ${B}$ and ${E}$ are the vectors of electric and magnetic field respectively as measured in the aforementioned inertial frame. $w(p,\rho)$ is the relativistic enthalpy per unit volume. In what follows, we use the equation of state
@@
where $\Gamma$ is the ratio of specific heats. Here we utilise the relativistic units where neither the speed of light no the geometric factor $1/4\pi$ appear in the equations (for example $\sigma=b^{2}/w$). We also agree that for any 3-vector of the space ${a}$, $a^{2}=a_{i}a^{i}$ and $a=\sqrt{a^{2}}$, and for any 4-vector $a^{\nu}$ of the spacetime, $a^{2}=a_{\nu}a^{\nu}$, and use $+$-$+$- + + + signature for the spacetime.
Let us now discuss how the errors in numerical integration of the energy-momentum equations can result in an unphysical state. Consider the 4-vector of energy-momentum density, $\Pi^{\mu}=-T^{\mu\nu}n_{\nu}$, where $T^{\mu\nu}$ is stress-energy-momentum tensor, and $n_{\nu}$ is the 4-velocity of the fiducial observer who measures the energy and momentum. When the observer is at rest in the space, $n_{\nu}=-\delta^{t}_{\nu}$ and $\Pi^{\nu}=({\cal E},\mn@boldsymbol{S})$, where ${\cal E}$ and ${S}$ are the energy and momentum densities respectively. For the electromagnetic field, this is
$\Pi_{\mbox{\tiny em}}^{\nu}=\left(\frac{E^{2}+B^{2}}{2},\mn@boldsymbol{E}\!\times\!\mn@boldsymbol{B}\right)\,,\quad\mbox{where}\quad\mn@boldsymbol{E}=-\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}\,,$ (8)
@@
$\Pi_{\mbox{\tiny em}}^{2}=-\frac{1}{4}\left(B_{\parallel}^{2}+\frac{B_{\perp}^{2}}{\gamma^{2}}\right)^{2}<0\,,$ (9)
@@
$\Pi^{\nu}_{\mbox{\tiny pl}}=(w\gamma^{2}-p,w\gamma^{2}\mn@boldsymbol{v})\,,$ (10)
@@
$\Pi_{\mbox{\tiny pl}}^{2}=\gamma^{2}w(2p-w)-p^{2}\,.$ (11)
@@
$\Pi_{\mbox{\tiny t}}^{2}=\Pi_{\mbox{\tiny em}}^{2}+\Pi_{\mbox{\tiny pl}}^{2}+2(-{\cal E}_{\mbox{\tiny em}}{\cal E}_{\mbox{\tiny pl}}+(\mn@boldsymbol{S}_{\mbox{\tiny em}}\!\cdot\!\mn@boldsymbol{S}_{\mbox{\tiny pl}}))<0$ (12)
and hence $\Pi^{\nu}_{\mbox{\tiny t}}$ is also time-like (This is a particular case of the general result on the sum of future-directed time-like vectors of spacetime.).